# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

### v2

@@ 3.7 Nonlinear Schröedinger Equation
Figure 19: Exact solution of the real part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 20: Best real-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 21: Exact solution of the imaginary part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 22: Best imaginary-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
@@
Table 6: Nonlinear Schroedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch Real part Imaginary part
$1.80\times 10^{-3}$ $2.58\times 10^{-3}$
$1.84\times 10^{-3}$ $3.27\times 10^{-3}$
$7.34\times 10^{-2}$ $1.23\times 10^{-1}$
$8.32\times 10^{-2}$ $1.39\times 10^{-1}$
$1.10\times 10^{-2}$ $2.10\times 10^{-2}$
$1.70\times 10^{-2}$ $3.27\times 10^{-2}$
@@
4 Conclusion
We provide an ablation study of hard-constraints and self-adaptive loss weights to study the solution dependency of time-dependent PDEs on its initial conditions. The transformation provided by hard-constraints significantly alleviates the training challenges typically associated with stiff time-dependent PDEs subject. Our study demonstrates effectiveness across seven distinct types of such PDEs, encompassing scalar-valued outputs, vector-valued outputs, and even complex-valued solutions. This approach not only enhances training efficiency but also improves the accuracy and stability of the learned solutions, particularly in the presence of stiffness, where conventional PINNs often struggle to converge.
Although we provide an ablation study of hard-constraints and self-adaptive loss weights, our list of possible training remedies for PINNs is not exhaustive. Moreover, our tests should also include elliptical PDEs with high oscillatory coefficients. Future work includes extending the hard-cosntraint framework to two- and three-dimensional PDEs, which pose additional challenges in both architecture design and computational cost. Ongoing research also focuses on identifying optimal transformations $\psi$ and $\phi$, which govern the embedding of IC and BC into the network structure, with the goal of further enhancing performance and generalizability.
Acknowledgements
BH and MZ thank Prof. Yannis Kevrekids (JHU) for the discussion of well-conditioning of PDE solution maps, and Prof. George Karniadakis (Brown) for his suggestion on training enhancement methods for PINNs. BH developed and implemented the algorithm, ran the simulation and analyzed the data. UB and MZ designed the reserach. All authors wrote the paper. MZ is supported by NSF-AoF grant #$\#2225507$ and Ralph E. Powe Junior Faculty Enhancement Award from ORAU for FY24242424.
Supplementary Material
The source code, initial training data, initial weights, Chebfun code, and ground truth solutions used for benchmarking will provided in https://github.com/baolihao/Training-PINNs-with-Hard-Constraints-and-Adaptive-Weights. The list of hyperparameters and computational resources used to produce the results will be discussed later in this appendix.

### v3

@@ B.6 Nonlinear Schröedinger Equation
Figure 9: Best real-part solutions of the Nonlinear Schröedinger Equation.
Figure 10: Best imaginary-part solutions of the Nonlinear Schröedinger Equation.
@@
Ablation Settings Performance: $\mathcal{E}_{2}$
PBC IC1 IC2 SA mini-batch Real part Imag part
$1.80\times 10^{-4}$ $2.58\times 10^{-4}$
$1.84\times 10^{-4}$ $3.27\times 10^{-4}$
$7.34\times 10^{-3}$ $1.23\times 10^{-2}$
$8.32\times 10^{-3}$ $1.39\times 10^{-2}$
$1.10\times 10^{-3}$ $2.10\times 10^{-3}$
$1.70\times 10^{-3}$ $3.27\times 10^{-3}$
@@
Table 6: Nonlinear Schröedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
B.7 Linear Advection
The linear advection equation models the transport of a quantity by a constant velocity field and serves as a fundamental benchmark for testing numerical methods due to its simplicity and rich structure LeVeque (2002); Toro (2013). It describes the propagation of a scalar profile without deformation and is given by:
$u_{t}+cu_{x}=0,\quad\quad\quad(t,x)\in(0,T]\times(0,2\pi),$ (11)
$u(0,x)=\sin(x),\quad x\in[0,2\pi],$
$u(t,0)=u(t,2\pi),\quad t\in[0,T].$
Unlike dissipative PDEs such as the Allen–Cahn or Cahn–Hilliard equations, the linear advection equation conserves the shape of the initial profile as it translates. This problem becomes “stiffer” as the advection velocity increases.
Theorem 3.
Let $u(t,x)$ denote the solution to the linear advection equation (11). With the transformed neural approximation in (3) and $\psi$ and $\eta$ satisfying the conditions in (4), then $u(t,x)$ satisfies the PDE iff $\eta$ satisfies the forced linear transport equation:
$\eta_{t}+c\eta_{x}=-(\psi_{t}+c\psi_{x}).$
This PDE is well-posed and stable for PINN training provided the following conditions hold:
  • The function $f(t,x):=-(\psi_{t}+c\psi_{x})$ satisfies $f\in C^{1}([0,T]\times\Omega)$.
  • If $f(t,x)$ is uniformly bounded or decays in time, then the solution $\eta$ remains stable and avoids growth-induced overfitting.
For our choice, we have $\psi(t,x)=u_{0}(x)e^{-\alpha t}$ and $\phi(t,x)=t$ with $\alpha>0$, ensuring $\eta(0,x)=0$. Then, the forcing term satisfies $\eta_{t}+c\eta_{x}=(\alpha\sin{x}-c\cos{x})e^{-\alpha t}$. Since this forcing is smooth and decays exponentially in time, it guarantees both the stability and regularity of the neural network $\eta(t,x)$. Here, we compare the results produced by our method with those from a baseline PINN under the same hyperparameter settings. In our experiments, both models were able to converge at low velocities; however, we observed that higher advection velocities introduced stiffness. Our method exhibited greater robustness and was less affected by stiffness compared to the baseline PINN.
Figure 11: Periodic advection results after 50,000 iterations.
The implementation details are the same as in Section B.1, except that we perform $10$ independent replications of the experiment to account for the randomness in neural network weight initialization, rather than using the same initial weights. Table 7 shows the averaged relative $L_{1}$$L_{2}$-error over $10$ independent repetitions. As the advection velocity increases and the PDE becomes stiffer, making it harder for baseline PINN to train. However, our method consistently produces better results. This trend is clearly reflected in Table 7. Even at low velocities, our method demonstrates significantly better accuracy compared to the baseline PINN. As $c$ increases, both the relative $L_{1}$ and $L_{2}$ errors grow, however, the errors associated with the PINN rise significantly faster than those of our method. The most striking example occurs at $c=50$, where the base line PINN fails to approximate the solution, while our method still produces stable and good results. This behavior is evident in Table 7 and Figure 11. Table 8 shows the ablation study of linear advection equation.
Solver $c$ Rel. $L_{1}$ error Rel. $L_{2}$ error
30 $0.0030$ $0.0041$
HC-PINN 40 $0.0141$ $0.0159$
50 $0.0147$ $0.0164$
60 $0.0171$ $0.0182$
30 $0.0478$ $0.0579$
CINN(Braga-Neto, 2022) 40 $0.0714$ $0.0852$
50 $0.4692$ $0.5365$
60 $0.7293$ $0.7134$
30 $0.0873$ $0.1003$
Baseline 40 $0.3918$ $0.4395$
PINN 50 $0.7140$ $0.7797$
60 $0.8339$ $0.8664$
Table 7: Comparison of HC-PINN, CINN and baseline PINN across different advection speeds $c$ over 10 independent repetitions.
Ablation Settings Performance
PBC IC SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.82\times 10^{-2}$ $1.03\times 10^{-2}$
$9.75\times 10^{-2}$ $8.65\times 10^{-2}$
$7.50\times 10^{-1}$ $7.87\times 10^{-1}$
$3.87\times 10^{-2}$ $5.47\times 10^{-2}$
$7.44\times 10^{-1}$ $8.01\times 10^{-1}$
$07\times 10^{-1} 1.1$$19\times 10^{-1}$
$65\times 10^{-1} 7.7$$71\times 10^{-1}$
$8.66\times 10^{-1}$ $8.34\times 10^{-1}$
Table 8: Linear Advection Equation (c=60): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.